Combustion reactions are a notable source of carbon dioxide in the environment. Using the following balanced equation, how many grams of carbon dioxide are formed when 100.00 g of propane (C3H8) is burned? Express your answer to the correct number of significant figures. Be sure to show all steps completed to arrive at the answer.


C3H8 +5O2 = 3CO2+4H2O



1
Expert's answer
2021-06-18T07:25:39-0400

C3H8 +5O2 = 3CO2+4H2O

According to the equation, n (CO2) = 3 x n (C3H8)

n = m / M

M (C3H8) = 44.1 g/mol

M (CO2) = 44.01 g/mol

n (C3H8) = 100.00 / 44.1 = 2.3 mol

n (CO2) = 3 x n (C3H8) = 3 x 2.3 = 6.9 mol

m (CO2) = 6.9 x 44.0 = 303.7 g